OpenStudy (diana):

How long does it take for an investment to double in value if it is invested at 8% per annum compounded monthly?

OpenStudy (anonymous):

solve \[2=(1+\frac{.08}{12})^{12t}\] for t

OpenStudy (anonymous):

first step is \[12t=\frac{\ln(2)}{\ln(1+\frac{.08}{12})}\] second step is \[ t=\frac{\ln(2)}{12\ln(1+\frac{.08}{12})}\] third step is use a calculator